the number of real values of x where the function f(x) = cosx+cos{(√2)x} attains its maximum is


(a) 0


(b) 1


(c) 2


(d) infinite

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear raman malik.

f(x) = cosx+cos{(√2)x}

f'(x)=-sin x -√2 sin √2x =0

    it gives x= 2m ∏/(√2 +1 )   and  (2m +1 ) ∏  /(√2 -1)  where m is any natural number

  clearly both values are not multiple of 2∏

only for m=0 

and f(0) = 1+1 =2 is its maximum value

so only one solution option b is correct

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin